OpenStudy (anonymous):

Which ordered pairs in the form (x, y) are solutions to the equation 3x - 4y = 21? Choose all answers that are correct. A. (-3, 3) B. (-1, -6) C. (7, 0) D. (11, 3)

OpenStudy (texaschic101):

3x - 4y = 21 lets check... A. (-3,3)...x = -3 and y = 3 now we sub 3(-3) - 4(3) = 21 -9 - 12 = 21 - 21 = 21 ---- not this one B. (-1,-6)...x = -1 and y = -6 now we sub 3(-1) - 4(-6) = 21 -3 + 24 = 21 21 = 21 (correct)---- this is one C. (7,0)...x = 7 and y = 0 now we sub 3(7) + 4(0) = 21 21 + 0 = 21 21 = 21 (correct) ---- this is one D. (11,3)...x = 11 and y = 3 3(11) + 4(3) = 21 33 + 12 = 21 45 = 21 --- not this one

OpenStudy (anonymous):

thank you so much,when you can.Can you figure out these 3 other questions?